OpenStudy (anonymous):

Find the distance between the points (0,15) and (8,0).

OpenStudy (anonymous):

sqrt(8^2 + 15^2 )

OpenStudy (anonymous):

sqrt(64 + 225) = 17

OpenStudy (anonymous):

answer is 17 units

OpenStudy (anonymous):

d = sqrt(15^2 + 8^2) =17

OpenStudy (anonymous):

More precisely, it is the square root of the change in each point. \[\sqrt{(x _{2 } - x_{1})^2 + (y_{2} - y_{1})^2}\] \[\sqrt{(8 - 0)^2 + (0 - 15)^2}\]

OpenStudy (anonymous):

misspac, do you understand how to use the distance formula? You've been asking a lot of very similar questions about it.

hero (hero):

There's a much easier way to understand this. If you understand pythagorean theorem, you can understand this